Economists frequently make use of quintiles (i.e., the 20th, 40th, 60th, and 80th percentiles) particularly

when discussing incomes. Suppose that in a large city household incomes are normally distributed

with a mean of $50,000 and a standard deviation of $10,000. An economist wishes to identify the

quintiles. Unfortunately, he did not pass his statistics course. Help him by providing the quintiles.


1
Expert's answer
2020-11-26T17:48:57-0500

Given :

mean= 50000; sd= 10000


To find 20th percentile, that is "p(Z_a)=0.2"


Using ti84, invnorm(0.2)= -0.842.Hence Z= -0.842.

therefore using Z score formula, "x=Z*sd+mean"

"x= -0.842*10000+50000=41580"


To find 40th percentile, that is "p(Z_a)=0.4"


Using ti84, invnorm(0.4)= -0.253.Hence Z= -0.253.

therefore using Z score formula,"x=Z*sd+mean\\"

"x= -0.253*10000+50000=47470"


To find 40th percentile, that is "p(Z_a)=0.6"


Using ti84, invnorm(0.6)= 0.253.Hence Z= 0.253.

therefore using Z score formula,"x=Z*sd+mean\\"

"x=0.253*10000+50000" =52530


To find 40th percentile, that is "p(Z_a)=0.8"


Using ti84, invnorm(0.8)=0.842.Hence Z= 0.842.

therefore using Z score formula,"x=Z*sd+mean\\"

"x= 0.842*10000+50000=58420"